OpenStudy (anonymous):

What is the slope of a line that contains the 2 points (10, 23) and (14, 28.6).

OpenStudy (anonymous):

Slope will be (28.6-23)/(14-10) = 5.6/4 = 1.4.

OpenStudy (anonymous):

thanks (::